# Spanning Training Progress: Temporal Dual-Depth Scoring (TDDS) for Enhanced Dataset Pruning

### v1

@@ Abstract
Dataset pruning aims to construct a coreset capable of achieving performance comparable to the original, full dataset. Most existing dataset pruning methods rely on snapshot-based criteria to identify representative samples, often resulting in poor generalization across various pruning and cross-architecture scenarios. Recent studies have addressed this issue by expanding the scope of training dynamics considered, including factors such as forgetting event and probability change, typically using an averaging approach. However, these works struggle to integrate a broader range of training dynamics without overlooking well-generalized samples, which may not be sufficiently highlighted in an averaging manner. In this study, we propose a novel dataset pruning method termed as Temporal Dual-Depth Scoring (TDDS), to tackle this problem. TDDS utilizes a dual-depth strategy to achieve a balance between incorporating extensive training dynamics and identifying representative samples for dataset pruning. In the first depth, we estimate the series of each sample’s individual contributions spanning the training progress, ensuring comprehensive integration of training dynamics. In the second depth, we focus on the variability of the sample-wise contributions identified in the first depth to highlight well-generalized samples. Extensive experiments conducted on CIFAR and ImageNet datasets verify the superiority of TDDS over previous SOTA methods. Specifically on CIFAR-100, our method achieves 54.51% accuracy with only 10% training data, surpassing random selection by 7.83% and other comparison methods by at least 12.69%.
@@ 1 Introduction
Explosively growing datasets [17, 50, 9] have been crucial for the success of deep neural networks (DNNs) in various applications [18, 48, 7, 32]. However, learning from large-scale datasets is both time-consuming and financially prohibitive [37, 36, 5, 24, 42, 46]. Actually, a substantial portion of data samples are redundant [41, 44, 49], which means they can be excluded from training without compromising performance.
@@
Numerous dataset pruning works [22, 29, 44, 47, 14, 19, 26] have explored various criteria [11, 38] for identifying important data samples, including distribution distance [43, 44], uncertainty [8], loss [41, 28], and gradient [28]. Most of them evaluate sample importance using a snapshot of training progress. For example, Paul et al. [28] uses the error vector score generated by a few epochs  [^1] Usually, the first $10$ or $20$ epochs. in early training; Xia et al. [44] calculates the distribution distances of features at the end of training.
@@
In this paper, we propose Temporal Dual-Depth Scoring (TDDS) pruning method, designed to effectively balance and address the aforementioned problems. We employ a dual-depth temporal scoring metric that spans the training process for data selection. In the inner level (depth one), we approximate the actual contribution of each sample in every epoch. This approximation is achieved by projecting the sample-wise gradients onto the accumulated gradient in each epoch. We then integrate these contributions over a specified window, typically spanning 10 epochs, to incorporate more training dynamics. In the outer level (depth two), we reduce reliance on the averaging down-sampling operation to prevent overlooking vital samples. Well-generalized samples are identified based on the variability of the projected gradients, calculated in each window from the inner level. This design of the outer level draws inspiration from He et al. [13], who tracked the variance of predicted target probabilities throughout training to enhance pruning effectiveness. The pipeline of our proposed TDDS is clearly illustrated in Figure 2. Extensive experiments demonstrate the superior generalization capability of the proposed TDDS, enabling it to outperform other comparison methods across different pruning settings on various datasets and networks. For example, on CIFAR-100, our method achieves 54.51% accuracy with only 10% training data, surpassing random selection by 7.83% and other SOTA methods by at least 12.69%.
@@ 2.1 Dataset Pruning
Dataset Pruning, also known as Coreset Selection aims to shrink the dataset scale by selecting important samples according to some predefined criteria. Herding [6, 43] and Moderate [44] calculate the distance in feature space. Entropy [8] and Cal [21] explore the uncertainty and decision boundary with the predicted outputs. GradND/EL2N [28] quantifies the importance of a sample with its gradient magnitude. Most of these methods are snapshot-based, that is they do not care about the training dynamics that actually distinguish the behaviors of samples in the overall optimization. There are some efficient training methods [40, 25, 45, 10], which focus more on training efficiency and execute online selection spanning training progress. For example, gradient matching-based methods GradMatch [15] and Craig [23] minimize the distance between gradients produced by the full dataset and coreset. ACS [14] selects samples with larger gradient magnitude to accelerate quantization-aware training. These methods can capture training dynamics but often necessitate repeated importance re-evaluating and can not be generalized to different networks. Some methods such as Forgetting [41] and AUM [30] also have proved training dynamics benefits samples importance assessment. However, the downsampling operation, which averages dynamics across training, ignores the variability which is crucial for evaluating sample generalization, ultimately leading to sub-optimal coresets.
@@ 2.2 Variance in Deep Learning
Variance, a standard second-order statistic, revealing the variability along a certain dimension, has been explored in numerous works [4, 27, 31]. For example, MVT [27] transfers the observed variance of one class to another to improve the robustness of the unseen examples and few-shot learning tasks. VICReg [4] maintains variance of embeddings to prevent the collapse in which two branches produce identical outputs regardless of inputs. For the gradient, its variance along the parameter dimension is an approximation of Hessian matrix indicating loss landscape. Based on that, Fishr [31] matches domain-level variances of gradients for out-of-distribution generalization. LCMat [34] induces a better adaptation of the pruned dataset on the perturbed parameter region than the exact point matching. Besides, calculating gradient variance along time dimension tracks dynamic optimization change. VoG [1] uses this to estimate sample importance and the visualization shows that high-variance samples are more worthwhile to learn. Dyn-Unc [13] calculates the time-dimension variance of predicted target class possibility to evaluate sample importance. Differing from VoG and Dyn-Unc, our TDDS does not need pixel-wise gradient calculation while considering non-target probabilities, which enables a more precise evaluation with lower computational overhead.
@@ 3 Preliminary
Throughout this paper, we denote the full training dataset as $\mathbb{U}=\left\{\left(\bm{x}_{n},\bm{y}_{n}\right)\right\}_{n=1}^{N}$, where $\bm{x}_{n}\in\mathbb{R}^{D}$ and $\bm{y}_{n}\in\mathbb{R}^{1\times C}$ are drawn i.i.d. from a natural distribution $\mathcal{D}$. We denote a neural network parameterized with weight vector $\bm{\theta}$ as $f_{\bm{\theta}}$. Under these definitions, the optimization objective of $f_{\bm{\theta}}$ on $\mathbb{U}$ is to minimize empirical risk $\mathcal{L}(\mathbb{U};\bm{\theta})=\frac{1}{N}\sum_{n=1}^{N}\ell\left(f_{\bm{\theta}}(\bm{x}_{n}),\bm{y}_{n}\right)$, where $f_{\bm{\theta}}(\bm{x}_{n})\in\mathbb{R}^{1\times C}$ outputs the predicted probabilities of each class. Typically, $\bm{\theta}$ is updated as follow  [^2] Here, $t$ indicates epoch and the batch-wise updates [2] in one epoch are omitted.,
@@ 4 Methodology
As discussed earlier, despite that exploring long-range dynamics provides a more comprehensive sample evaluation, the averaging operation hinders the construction of a coreset with strong generalization. In this section, we present the proposed Temporal Dual-Depth Scoring strategy, which balances the incorporation of training dynamics and the identification of well-generalized samples. In the outer level, we identify samples with their variability during training which is encouraged theoretically by the optimization space matching between coreset and original dataset (refer to Section 4.1). In the inner level, the gradient in the outer level is replaced with the projected component on the accumulated direction to estimate valid contributions in every epoch (refer to Section 4.2). Finally, to enhance efficiency, we conduct TDDS with a moving window (refer to Section 4.3). The proposed TDDS is summarized in Algorithm 1.
@@ 4.1 Outer Level: Enhance Coreset Generalization by Maximizing Variance
To construct the coreset $\mathbb{S}$ satisfying Equation 2, gradient-based methods [15, 23, 3] propose to minimize the distance between coreset and full dataset gradient,
$\mathbb{S}^{*}=\operatorname*{arg\,min}_{\mathbb{S}\subset\mathbb{U}}$ $\|\bm{G}_{t,\mathbb{U}}-\bm{G}_{t,\mathbb{S}}\|,\quad\mbox{where}$ (3)
$\bm{G}_{t,\mathbb{U}}=\sum_{\begin{subarray}{c}n=1,\bm{x}_{n}\in\mathbb{U}\end{subarray}}^{|\mathbb{U}|}\bm{g}_{t}(\bm{x}_{n}),\quad$ $\bm{G}_{t,\mathbb{S}}=\sum_{\begin{subarray}{c}m=1,\bm{x}_{m}\in\mathbb{S}\end{subarray}}^{|\mathbb{S}|}\bm{g}_{t}(\bm{x}_{m}).$
To solve this minimization problem, [14] recommends pruning samples with smaller gradient magnitudes to minimize their potential impact on optimization.
However, whenever the distance is minimized, the obtained snapshot-based coreset fails to generalized across training progress. Here are the reasons: (1) the distinctive contributions of samples are temporal statistics. Evaluating them at a specific time point causes significant bias when considering dynamic optimization; (2) samples with larger individual gradients dominate the matching process. Without considering the current optimization context, these samples introduce a substantial error in approximating the accumulated gradient.
To address the first issue, we proposed to extend Equation 3 to a continuous training duration to capture training dynamics,
$\mathbb{S}^{*}$ $=\operatorname*{arg\,min}_{\mathbb{S}\subset\mathbb{U}}\frac{1}{T}\sum_{t=1}^{T}\|\bm{\mathcal{G}}_{t,\mathbb{U}}-\bm{\mathcal{G}}_{t,\mathbb{S}}\|^{2},$ (4)
Here, the gradients generated by $\mathbb{S}$ and $\mathbb{U}$ at time $t$ are defined as $\bm{\mathcal{G}}_{t,\mathbb{U}}=[|\bm{g}_{t}(\bm{x}_{n})|]_{n=1}^{N}$ and $\bm{\mathcal{G}}_{t,\mathbb{S}}=[|\bm{g}_{t}(\bm{x}_{m})|]_{m=1}^{M}$, respectively. Experiments in Section 5.4 shows that magnitude-based minimization performs better than gradient-based when considering continuous training.
@@
$\mathbb{S}^{*}=\operatorname*{arg\,max}_{\mathbb{S}\subset\mathbb{U}}\mathcal{R}(\mathbb{S}),\quad\mathcal{R}(\mathbb{S})=\sum_{t=1}^{T}\|\bm{\mathcal{G}}_{t,\mathbb{S}}-\overline{\bm{\mathcal{G}}}_{\mathbb{S}}\|^{2}.$ (5)
Here $\overline{\bm{\mathcal{G}}}_{\mathbb{S}}=\frac{1}{T}\sum_{t=1}^{T}\bm{\mathcal{G}}_{t,\mathbb{S}}$ is the gradient mean over training time. Detailed proof can be founded in Supplementary. Equation 5 measures the variability of samples and enhances the identification of well-generalized samples compared to previous average-based methods [41, 30].
@@ 4.2 Inner Level: Estimate Actual Contribution with Projected Gradient
where $\overline{|\bm{g}(\bm{x}_{n})|}=\frac{1}{T}\sum_{t=1}^{T}|\bm{g}_{t}(\bm{x}_{n})|$. Equation 6 treats sample-wise gradient independently, which may not precisely describe the valid contribution in the current optimization context. That is the second problem mentioned in Section 4.1. As illustrated in Figure 3(a), the gradient of sample $\bm{x}_{n}$ remains constant between $t$ and $t+1$. However, when incorporating it into the current optimization context, its valid contribution undergoes a significant shift. Thus, we propose to calculate the gradient variance after projecting it on the accumulated direction,
$|\bm{g}_{t}(\bm{x}_{n})|$ $=|\bm{g}_{t}(\bm{x}_{n})\cdot\sum_{\begin{subarray}{c}n=1,\bm{x}_{n}\in\mathbb{U}\end{subarray}}^{|\mathbb{U}|}\bm{g}_{t}(\bm{x}_{n})|.$ (7)
@@
Considering that $f_{\bm{\theta}_{t}(\bm{x}_{n})}$ follows a Bernoulli distribution $\mathcal{B}(p_{t})$, where $p_{t}$ represents the probability of the target class at time $t$. If $p_{t}$ equals $p_{t+1}$, Equation 9 fails to distinguish, but Equation 10 can still differentiate the contributions of samples with $q_{t}$ and $q_{t+1}$. Here, $q_{t}$ and $q_{t+1}$ are the non-target probability distributions. This effectiveness has been proved in Section 5.4.

### v3

@@ Abstract
Dataset pruning aims to construct a coreset capable of achieving performance comparable to the original, full dataset. Most existing dataset pruning methods rely on snapshot-based criteria to identify representative samples, often resulting in poor generalization across various pruning and cross-architecture scenarios. Recent studies have addressed this issue by expanding the scope of training dynamics considered, including factors such as forgetting event and probability change, typically using an averaging approach. However, these works struggle to integrate a broader range of training dynamics without overlooking well-generalized samples, which may not be sufficiently highlighted in an averaging manner. In this study, we propose a novel dataset pruning method termed as Temporal Dual-Depth Scoring (TDDS), to tackle this problem. TDDS utilizes a dual-depth strategy to achieve a balance between incorporating extensive training dynamics and identifying representative samples for dataset pruning. In the first depth, we estimate the series of each sample’s individual contributions spanning the training progress, ensuring comprehensive integration of training dynamics. In the second depth, we focus on the variability of the sample-wise contributions identified in the first depth to highlight well-generalized samples. Extensive experiments conducted on CIFAR and ImageNet datasets verify the superiority of TDDS over previous SOTA methods. Specifically on CIFAR-100, our method achieves 54.51% accuracy with only 10% training data, surpassing baselines methods by more than 12.69%. Our codes are available at https://github.com/zhangxin-xd/Dataset-Pruning-TDDS.
@@ 1 Introduction
Explosively growing datasets [21, 53, 9, 54] have been crucial for the success of deep neural networks (DNNs) in various applications [22, 51, 7, 36]. However, learning from large-scale datasets is both time-consuming and financially prohibitive [41, 40, 5, 28, 45, 49]. Actually, a substantial portion of data samples are redundant [44, 47, 52], which means they can be excluded from training without compromising performance.
@@
Numerous dataset pruning works [26, 33, 47, 50, 17, 23, 30, 16, 12] have explored various criteria [13, 42] for identifying important data samples, including distribution distance [46, 47], uncertainty [8], loss [44, 32], and gradient [32]. Most of them evaluate sample importance using a snapshot of training progress. For example, Paul et al. [32] uses the error vector score generated by a few epochs [^1] Usually, the first $10$ or $20$ epochs. in early training; Xia et al. [47] calculates the distribution distances of features at the end of training.
@@
In this paper, we propose Temporal Dual-Depth Scoring (TDDS) pruning method, designed to effectively balance and address the aforementioned problems. We employ a dual-depth temporal scoring metric that spans the training process for data selection. In the inner level (depth one), we approximate the actual contribution of each sample in every epoch. This approximation is achieved by projecting the sample-wise gradients onto the accumulated gradient in each epoch, following the approach suggested by Du et al. [11]. We then integrate these contributions over a specified window, typically spanning 10 epochs, to incorporate more training dynamics. In the outer level (depth two), we reduce reliance on the averaging down-sampling operation to prevent overlooking vital samples. Well-generalized samples are identified based on the variability of the projected gradients, calculated in each window from the inner level. This design of the outer level draws inspiration from He et al. [15], who tracked the variance of predicted target probabilities throughout training to enhance pruning effectiveness. The pipeline of our proposed TDDS is clearly illustrated in Figure 2. Extensive experiments demonstrate the superior generalization capability of the proposed TDDS, enabling it to outperform other comparison methods across different pruning settings on various datasets and networks. For example, on CIFAR-100, our method achieves 54.51% accuracy with only 10% training data, surpassing random selection by 7.83% and other SOTA methods by at least 12.69%.
@@ 2.1 Dataset Pruning
Dataset Pruning, also known as Coreset Selection aims to shrink the dataset scale by selecting important samples according to some predefined criteria. Herding [6, 46] and Moderate [47] calculate the distance in feature space. Entropy [8] and Cal [25] explore the uncertainty and decision boundary with the predicted outputs. GradND/EL2N [32] quantifies the importance of a sample with its gradient magnitude. Most of these methods are snapshot-based, that is they do not care about the training dynamics that actually distinguish the behaviors of samples in the overall optimization. There are some efficient training methods [43, 29, 48, 10], which focus more on training efficiency and execute online selection spanning training progress. For example, gradient matching-based methods GradMatch [19] and Craig [27] minimize the distance between gradients produced by the full dataset and coreset. ACS [17] selects samples with larger gradient magnitude to accelerate quantization-aware training. These methods can capture training dynamics but often necessitate repeated importance re-evaluating and can not be generalized to different networks. Some methods such as Forgetting [44] and AUM [34] also have proved training dynamics benefits samples importance assessment. However, the downsampling operation, which averages dynamics across training, disregards the variability crucial for evaluating sample generalization, ultimately resulting in suboptimal coresets.
@@ 2.2 Variance in Deep Learning
Variance, a standard second-order statistic, revealing the variability along a certain dimension, has been explored in numerous works [4, 31, 35]. For example, MVT [31] transfers the observed variance of one class to another to improve the robustness of the unseen examples and few-shot learning tasks. VICReg [4] maintains variance of embeddings to prevent the collapse in which two branches produce identical outputs regardless of inputs. For the gradient, its variance along the parameter dimension is an approximation of Hessian matrix indicating loss landscape. Based on that, Fishr [35] matches domain-level variances of gradients for out-of-distribution generalization. LCMat [38] induces a better adaptation of the pruned dataset on the perturbed parameter region than the exact point matching. Besides, calculating gradient variance along time dimension tracks dynamic optimization change. VoG [1] uses this to estimate sample importance and the visualization shows that high-variance samples are more worthwhile to learn. Dyn-Unc [15] calculates the time-dimension variance of predicted target class probability to evaluate sample importance. Differing from VoG and Dyn-Unc, our TDDS does not need pixel-wise gradient calculation while considering non-target probabilities, which enables a more precise evaluation with lower computational overhead.
@@ 3 Preliminary
Throughout this paper, we denote the full training dataset as $\mathbb{U}=\left\{\left(\bm{x}_{n},\bm{y}_{n}\right)\right\}_{n=1}^{N}$, where $\bm{x}_{n}\in\mathbb{R}^{D}$ and $\bm{y}_{n}\in\mathbb{R}^{1\times C}$ are drawn i.i.d. from a natural distribution $\mathcal{D}$. We denote a neural network parameterized with weight matrix $\bm{\theta}$ as $f_{\bm{\theta}}$. Under these definitions, the optimization objective of $f_{\bm{\theta}}$ on $\mathbb{U}$ is to minimize empirical risk $\mathcal{L}(\mathbb{U};\bm{\theta})=\frac{1}{N}\sum_{n=1}^{N}\ell\left(f_{\bm{\theta}}(\bm{x}_{n}),\bm{y}_{n}\right)$, where $f_{\bm{\theta}}(\bm{x}_{n})\in\mathbb{R}^{1\times C}$ outputs the predicted probabilities of each class. Typically, $\bm{\theta}$ is updated as follow [^2] Here, $t$ indicates epoch and the batch-wise updates [2] in one epoch are omitted.,
@@ 4 Methodology
As discussed earlier, despite that exploring long-range dynamics provides a more comprehensive sample evaluation, the averaging operation hinders the construction of a coreset with strong generalization. In this section, we present the proposed Temporal Dual-Depth Scoring (TDDS) strategy, which balances the incorporation of training dynamics and the identification of well-generalized samples. In the outer level, we identify samples with their variability during training which is encouraged theoretically by the optimization space matching between coreset and original dataset (refer to Section 4.1). In the inner level, the gradient in the outer level is replaced with the projected component on the accumulated direction to estimate valid contributions in every epoch (refer to Section 4.2). Finally, to enhance efficiency, we conduct TDDS with a moving window (refer to Section 4.3). The proposed TDDS is summarized in Algorithm 1.
@@ 4.1 Outer Level: Enhance Coreset Generalization by Maximizing Variance
To construct the coreset $\mathbb{S}$ satisfying Equation 2, gradient-based methods [19, 27, 3] propose to minimize the distance between full dataset and coreset gradient,
$\mathbb{S}^{*}=\operatorname*{arg\,min}_{\mathbb{S}\subset\mathbb{U}}$ $\|\bm{G}_{t,\mathbb{U}}-\tilde{\bm{G}}_{t,\mathbb{S}}\|,\quad\mbox{where}$
$\bm{G}_{t,\mathbb{U}}=\sum_{\begin{subarray}{c}n=1,\\ \bm{x}_{n}\in\mathbb{U}\end{subarray}}^{|\mathbb{U}|}\bm{g}_{t}(\bm{x}_{n}),\quad$ $\tilde{\bm{G}}_{t,\mathbb{S}}=\sum_{\begin{subarray}{c}m=1,\\ \bm{x}_{m}\in\mathbb{S}\end{subarray}}^{|\mathbb{S}|}\bm{g}_{t}(\bm{x}_{m}),$ (3)
$\tilde{\bm{G}}_{t,\mathbb{S}}$ can be regarded as the gradient reconstruction in coreset space. To solve this minimization problem, Huang et al. [17] prunes samples with smaller gradient magnitudes to minimize their potential impact on optimization.
However, whenever the distance is minimized, the obtained snapshot-based coreset fails to generalize across training progress. Here are the reasons: (1) the distinctive contributions of samples are temporal statistics. Evaluating them at a specific time point causes significant bias when considering dynamic optimization; (2) samples with larger individual gradients dominate the matching process. Without considering the current optimization context, these samples introduce a substantial error in approximating the accumulated gradient.
To address the first issue, we proposed to extend subsection 4.1 to a continuous training duration to capture training dynamics,
$\mathbb{S}^{*}$ $=\operatorname*{arg\,min}_{\mathbb{S}\subset\mathbb{U}}\frac{1}{T}\sum_{t=1}^{T}\|{\bm{\mathcal{G}}}_{t,\mathbb{U}}-\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}\|^{2},$ (4)
Here, the gradients generated by $\mathbb{U}$ at epoch $t$ are defined as $\bm{\mathcal{G}}_{t,\mathbb{U}}=[|\bm{g}_{t}(\bm{x}_{n})|]_{n=1}^{N}$ and similarly $\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}$ is its reconstructed counterpart using $\mathbb{S}$. Here we use magnitude instead of gradient itself to balance the training dynamic integration and complexity. Experiments in Section 5.4 also shows that magnitude-based minimization performs better than gradient-based when considering continuous training.
@@
$\mathbb{S}^{*}=\operatorname*{arg\,max}_{\mathbb{S}\subset\mathbb{U}}\mathcal{R}(\mathbb{S}),\quad\mathcal{R}(\mathbb{S})=\sum_{t=1}^{T}\|\bm{\mathcal{G}}_{t,\mathbb{S}}-\overline{\bm{\mathcal{G}}}_{\mathbb{S}}\|^{2},$ (5)
where $\overline{\bm{\mathcal{G}}}_{\mathbb{S}}=\frac{1}{T}\sum_{t=1}^{T}\bm{\mathcal{G}}_{t,\mathbb{S}}$, $\bm{\mathcal{G}}_{t,\mathbb{S}}=[|\bm{g}_{t}(\bm{x}_{m})|]_{m=1}^{M}$ is the mean gradient over training time. Detailed proof can be founded in Supplementary. Equation 5 measures the variability of samples and enhances the identification of well-generalized samples compared to previous average-based methods [44, 34].
@@ 4.2 Inner Level: Estimate Actual Contribution with Projected Gradient
where $\overline{|\bm{g}(\bm{x}_{n})|}=\frac{1}{T}\sum_{t=1}^{T}|\bm{g}_{t}(\bm{x}_{n})|$. Equation 6 treats sample-wise gradient independently, which may not precisely describe the valid contribution in the current optimization context. That is the second problem mentioned in Section 4.1. As illustrated in Figure 3(a), the gradient of sample $\bm{x}_{n}$ remains constant between $t$ and $t+1$. However, when incorporating it into the current optimization context, its valid contribution undergoes a significant shift. Thus, we propose to calculate the gradient variance after projecting it on the accumulated direction, that is, $|\bm{g}_{t}(\bm{x}_{n})|$ is replaced by,
$|\bm{g}_{t}(\bm{x}_{n})|$ $\Leftarrow|\bm{g}_{t}(\bm{x}_{n})\cdot\sum_{\begin{subarray}{c}j=1,\bm{x}_{j}\in\mathbb{U}\end{subarray}}^{|\mathbb{U}|}\bm{g}_{t}(\bm{x}_{j})|.$ (7)
@@
Considering that $f_{\bm{\theta}_{t}(\bm{x}_{n})}$ follows a Bernoulli distribution $\mathcal{B}(p_{t})$, where $p_{t}$ represents the probability of the target class at epoch $t$. If $p_{t}$ equals $p_{t+1}$, Equation 9 fails to distinguish, but Equation 10 can still differentiate the contributions of samples with $q_{t}$ and $q_{t+1}$. Here, $q_{t}$ and $q_{t+1}$ are the non-target probability distributions. This effectiveness has been proved in Section 5.4.